Solve loga(7x+1)=loga(4x+16)
Answer by stanbon(48535)

Solve loga(7x+1)=loga(4x+16)
Take the anti-log of both sides to get:
7x+1 = 4x+16
3x=15
x=5
============
Cheers,
Stan H.